Sheridan Corp. is expected to grow rapidly at a rate of 35 percent for the next seven years. The company's first dividend, to be paid three years from now, will be $5. After seven years, the company (and the dividends it pays) will grow at a rate of 7.6 percent. What is the value of Sheridan stock with a required rate of return of 14 percent? (Round intermediate calculations and final answer to 2 decimal places, e.g. 15.20.)
Answer #1

Given:

D0= 5

Growth rate ( 1 to 7) = 35%

Growth rate (after 7) = 7.6%

D1= D0 (1+g) = 5(1+0.35) = 6.75

D2= D1 (1+g) = 6.75(1+0.35)= 9.1125

D3= D2 (1+g) = 9.1125(1+0.35) = 12.30

D4= D3 (1+g) = 12.30(1+0.35) = 16.605

D5= D4 (1+g) = 16.605 (1+0.076) = 17.87

Value = D/(1+r)^3 + D1/(1+r)^4 + D2/(1+r)^5 + D3/(1+r)^6 + D4/(1+r)^7 + (D5/ R - g)/(1+r)^7

= 5/(1+0.14)^3 + 6.75/(1+0.14)^4 + 9.1125/(1+0.14)^5 + 12.30/(1+0.14)^6 + 16.605/(1+0.14)^7 + (17.87/0.14 - 0.076) / (1+0.14)^7

= 5/(1.14)^3 + 6.75/(1.14)^4 + 9.1125/(1.14)^5 + 12.30/(1.14)^6 + 16.605/(1.14)^7 + (17.87/0.064) / (1.14)^7

= 5/1.4815 + 6.75/1.6889 + 9.1125/1.9254 + 12.30/2.195 + 16.605/ 2.50 + 279.21875 / 2.50

= 3.37 + 4 + 4.73 + 5.60 + 6.642 + 111.6875

= $136.03